# High-dimensional limit theorems for SGD: Momentum and Adaptive Step-sizes

### v1

@@ 4.5. Auxiliary $L^{2}$-Bound
We establish a technical lemma essential for controlling the accumulation of second-order stochastic errors in Lemma 4.8,4.6, 4.7. We recall the following random variables representing the stochastic interactions:
@@ 5. Proofs for Tensor PCA
The first part of the proof pertaining to SGD-M follows from Theorem 2.3 and Section 7 of [benarous2024high] along with the verification of the final item of $\delta_{n}$-localizability (which is verified by the same bounds as the previous item). Moving on to SGD-U, we split the proof into two parts. The proof also follows from Theorem 2.3 along with the discussion of section 2.1. Thus we simply must verify asymptotic closability and $\delta_{n}$-localizability. We prove these two assumptions separately below starting with asymptotic closability.
@@ 5.1.1. Proof of Asymptotic Closability
We check asymptotic closability starting with the pre-limits for $\nabla\tilde{\Phi}$ and $\tilde{V}$ under the normalized gradient pre-conditioner. Recall that from [benarous2024high, Section 7], $\nabla L$ is a Gaussian vector with mean $\nabla\Phi$ and covariance $V$ given by
@@
$\mathbb{E}\left[\frac{\nabla L}{\|\nabla L\|}\right]$ $=\frac{\nabla\Phi}{D}-\frac{1}{2D^{3}}O(1)+O(n^{-2})=\frac{\nabla\Phi}{D}+O(n^{-3/2})$
@@
which is $O(n^{-2})$ again by Lemma 7.4. For the leading two terms, we use Stein’s lemma for vector-valued functions (e.g. see [liu2016kernelized, Lemma 2.2]),
@@
5.1.2. Proof of $\delta_{n}$-localizability
@@
We verify $\delta_{n}$-localizability for tensor pca with $u(x)=\sqrt{n}(m(x),r_{\perp}^{2}(x))$. We note that verification will then follow immediately for the cases $u(x)=(\sqrt{n}m(x),r_{\perp}^{2}(x))$ and $u(x)=(m(x),r_{\perp}^{2}(x))$. We also note item $(1)$ of $\delta_{n}$-localizability has already been verified for the no-preconditioning case. We move on to $(2)$:
$\sup_{x\in u^{-1}(E_{k})}\sqrt{n}\nabla\tilde{\Phi}$ $=\sup_{x\in u^{-1}(E_{k})}\sqrt{n}\frac{\|\nabla\Phi(x)\|}{\sqrt{\mathbb{E}\|\nabla L(x,y)\|^{2}}}+O(1/\sqrt{n})$
@@
$\sup_{x\in u^{-1}(E_{k})}\sqrt{n}\nabla\tilde{\Phi}$ $\lesssim\sup_{x\in u^{-1}(E_{k})}\sqrt{n}[\frac{C_{k}[\|x\|^{2k-2}+\|x\|^{k-1}]}{\sqrt{n\|x\|^{2k-2}+\|\Phi(x)\|^{2}}}+O(1/\sqrt{n})]\leq C_{k}$
@@
$\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\|\sqrt{n}\nabla\tilde{H}(x,y)\|^{8}$ $\lesssim n^{4}\sup_{x\in u^{-1}(E_{k})}\mathbb{E}[\|\frac{\nabla L(x,y)}{\|\nabla L(x,y)\|}\|^{8}+\|\nabla\tilde{\Phi}(x)]\|^{8}+O(1/n^{4})]$
@@
$\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\langle\sqrt{n}\nabla\tilde{H}(x,y),\nabla u_{i}(x)\rangle^{4}$ $=n^{2}\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\langle\frac{\nabla L(x,y)}{\|\nabla L(x,y)\|}-\nabla\tilde{\Phi}(x),\nabla u_{i}(x)\rangle^{4}$
$\lesssim n^{2}\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\langle\frac{\nabla L(x,y)}{\|\nabla L(x,y)\|},\nabla u_{i}(x)\rangle^{4}$
$\leq n^{2}\sup_{x\in u^{-1}(E_{k})}\sqrt{\mathbb{E}\langle\nabla L(x,y),\nabla u_{i}(x)\rangle^{8}\mathbb{E}\frac{1}{\|\nabla L(x,y)\|^{8}}}$
Where in the second line, we use $(a+b)^{4}\lesssim a^{4}+b^{4}$ and Jensen’s Inequality to remove the $\nabla\tilde{\Phi}$ term at the cost of a constant factor. We now recall that $\nabla L(x,y)$ is n-dimensional Gaussian with mean $O(1)$ and covariance $\Sigma=I_{d}+\lambda xx^{t},\|\Sigma\|_{op}=O(1),\|\Sigma\|_{F}=\Theta(n)$. For both $\sqrt{n}m$ and $\sqrt{n}r_{\perp}^{2}$ we have that $\langle\nabla L(x,y),\nabla u_{i}(x)\rangle$ is gaussian with variance bounded by $C_{k}n$ and by a technical lemma for Gaussians we have that $\mathbb{E}\frac{1}{\|\nabla L(x,y)\|^{8}}=O(1/n^{4})$ yielding:
$\max_{i}\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\langle\sqrt{n}\nabla\tilde{H}(x,y),\nabla u_{i}(x)\rangle^{4}=O(n^{2})$
@@
$\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\langle\nabla^{2}u_{i}(x),n\nabla\tilde{H}(x,y)\otimes\nabla\tilde{H}(x,y)-\tilde{V}\,\rangle^{2}$
$=n^{3}\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\,\text{tr}([I-vv^{t}](\nabla\tilde{H}(x,y)\otimes\nabla\tilde{H}(x,y)-\tilde{V}))^{2}$
$\lesssim n^{3}\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\,\text{tr}(\nabla\tilde{H}(x,y)\otimes\nabla\tilde{H}(x,y)-\tilde{V})^{2}+\mathbb{E}\text{tr}(vv^{t}[\nabla\tilde{H}(x,y)\otimes\nabla\tilde{H}(x,y)-\tilde{V}])^{2}$
We now bound each of the two terms separately starting with the first:
$\text{tr}(\nabla\tilde{H}(x,y)\otimes\nabla$ $\tilde{H}(x,y)-\tilde{V})=\text{tr}([\frac{\nabla L(x,y)}{\|\nabla L(x,y)\|}-\nabla\tilde{\Phi}(x)]^{\otimes 2}-\mathbb{E}[\frac{\nabla L(x,y)}{\|\nabla L(x,y)\|}-\nabla\tilde{\Phi}]^{\otimes 2})$
$=\text{tr}(\frac{\nabla L(x,y)^{\otimes 2}}{\|\nabla L(x,y)\|^{2}})-2\text{tr}(\frac{\nabla L(x,y)}{\|\nabla L(x,y)\|}\otimes\nabla\tilde{\Phi}(x))+\text{tr}(\nabla\tilde{\Phi}(x)^{\otimes 2})$
$-\mathbb{E}[\text{tr}(\frac{\nabla L(x,y)^{\otimes 2}}{\|\nabla L(x,y)\|^{2}})-2\text{tr}(\frac{\nabla L(x,y)}{\|\nabla L(x,y)\|}\otimes\nabla\tilde{\Phi}(x))+\text{tr}(\nabla\tilde{\Phi}(x)^{\otimes 2})]$
$=2\text{tr}(\frac{\nabla L(x,y)}{\|\nabla L(x,y)\|}\otimes\nabla\tilde{\Phi}(x))-\mathbb{E}2\text{tr}(\frac{\nabla L(x,y)}{\|\nabla L(x,y)\|}\otimes\nabla\tilde{\Phi}(x))$
@@
$\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\text{tr}(\nabla\tilde{H}(x,y)\otimes\nabla\tilde{H}(x,y)-\tilde{V})^{2}$ $\lesssim\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\text{tr}(\frac{\nabla L(x,y)}{\|\nabla L(x,y)\|}\otimes\nabla\tilde{\Phi}(x))^{2}$
$\leq\sup_{x\in u^{-1}(E_{k})}\|\nabla\tilde{\Phi}(x)\|^{2}$
$\leq C_{k}/n$
Where in the first line we use the fact $(a-b)^{2}\lesssim a^{2}+b^{2}$ and Jensen’s inequality. In the second line we use Cauchy Schwartz to bound the trace (which is here equal to an inner product) and in the last line we use item (2) of $\delta_{n}$-localizability above. Moving on to the second term, notice that by the first part of item (3) of $\delta_{n}$-localizability we have that:
$\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\text{tr}(vv^{t}[\nabla\tilde{H}(x,y)\otimes\nabla\tilde{H}(x,y)-\tilde{V}])^{2}$ $\lesssim\sup_{x\in u^{-1}(E_{k})}\text{tr}(vv^{t}[\nabla\tilde{H}(x,y)\otimes\nabla\tilde{H}(x,y))^{2}$
$=\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\langle\nabla\tilde{H}(x,y),v\rangle^{4}\leq C_{k}/n^{2}$
Which yields the desired:
@@
The last part of item $(3)$ is in fact not necessary as we consider $\beta=0$ for SGD-U. However it can easily be shown following the same steps as above and verifying that $\text{tr}(\frac{\nabla L(x_{1},y_{1})}{\|\nabla L(x_{1},y_{1})\|}\otimes\frac{\nabla L(x_{2},y_{2})}{\|\nabla L(x_{2},y_{2})\|})=O(1/\sqrt{n})$, which follows from computing the inner product of independent gaussian vectors and some basic bounds. ∎
@@
Following Theorem 2.3 and the discussion following it, we find that the dynamics of SGD-M follow those of online SGD up to the additional $\beta$-dependent constants. With this, our analysis immediately follows from [benarous2024high, Section 7] where we take the mappings,
@@
for $\beta\in(0,1)$. For example, by [benarous2024high, Proposition 3.1], the limiting generators for online SGD are given by
@@
The proof of Proposition 3.2 follows directly from the case of online SGD in [benarous2024high] and Theorem 2.3. In particular the equivalence of SGD and SGD-M after appropriate time and step-size rescaling. Below we prove Proposition 3.3.
@@
Applying Theorem 2.3 in light of 2.1, we simply check asymptotic closability, noting that $\delta_{n}$-localizability was already verified for this setting in the proof of Proposition 3.1. We check asymptotic closability which also follows easily from the same proof. The drifts under the new coordinates are written as
@@
$\tilde{f}_{\tilde{u}_{1}}=\begin{cases}\tilde{u}_{2}^{k-1}\tilde{u}_{1}-\lambda\left(\dfrac{\tilde{u}_{1}}{\sqrt{\tilde{u}_{2}}}\right)^{k-1}&k=2\\ \tilde{u}_{2}^{k-1}\tilde{u}_{1}&k\geq 3\end{cases},\quad\quad\tilde{f}_{\tilde{u}_{2}}=2\tilde{u}_{2}^{(k+1)/2}$
@@ 6. Proofs for Single-Index Models
We start with SGD-M below. The limiting generators were first determined in [single_index, Theorem 3.1, Corollary 3.2]. We provide a slightly different derivation of the result for completeness. Refer to [single_index, Lemma 3.5] to show that the problem satisfies the $\delta$-localizability and -closability assumptions.
@@
We now prove the remainder of the proposition for SGD-U. We note that since $\nabla L$ is a Gaussian vector with positive definite covariance structure as in the spiked tensor model, verifying $\delta$-localizability will be nearly identical using the same arguments and controls. With this, we focus on establishing closability.
Proof for SGD-U Dynamics.
@@
where $S^{2}=1$ almost surely. Since $a$ is an isotropic Gaussian vector, $\hat{a}$ is uniformly distributed on the sphere. It follows that $\mathbb{E}[\hat{a}\hat{a}^{\top}]=I/n$ (see, for instance [vershynin2018high, Proposition 3.3.8] for the construction). It follows then that the covariance matrix is:

### v2

@@ 6.5. Auxiliary $L^{2}$-Bound
We establish a technical lemma essential for controlling the accumulation of second-order stochastic errors in Lemma 6.8,6.6, 6.7. We recall the following random variables representing the stochastic interactions:
@@ 7. Proofs for Tensor PCA
The first part of the proof pertaining to SGD-M follows from Theorem 2.3 and Section 7 of [5] along with the verification of the final item of (3) in $\delta_{n}$-localizability (which is verified by the same bounds as the previous item). For example,
$\sup_{x\in E^{*}_{K,n}}\mathbb{E}_{y_{1}\perp y_{2}}\left[\langle\nabla^{2}u(x),\nabla H(x,y_{1})\otimes\nabla H(x,y_{2})\rangle^{2}\right]\lesssim\sup_{x\in E^{*}_{K,n}}\mathbb{E}\|\nabla H(x)\|^{4}\lesssim n^{2}$
for $u=(m,r^{2})$. For our diffusive limits about $m=0$, we have $\nabla^{2}u=0$ so that the bound is trivial. Moving on to SGD-U, we split the proof into two parts. The proof follows from the preconditioning theorem provided in appendix 5.3. Thus we simply must verify asymptotic closability and $(\delta_{n},\eta_{n})$-localizability. We prove these two assumptions separately below starting with asymptotic closability.
@@ 7.1.1. Proof of Asymptotic Closability
We check asymptotic closability starting with the pre-limits for $\nabla\tilde{\Phi}$ and $\tilde{V}$ under the normalized gradient pre-conditioner. Recall that from [5, Section 7], $\nabla L$ is a Gaussian vector with mean $\nabla\Phi$ and covariance $V$ given by
@@
$\mathbb{E}\left[\frac{\nabla L}{\|\nabla L\|}\right]$ $=\frac{\nabla\Phi}{D}-\frac{1}{2D^{3}}O(1)+O(d^{-2})=\frac{\nabla\Phi}{D}+O(n^{-3/2})$
@@
which is $O(n^{-2})$ again by Lemma 9.4. For the leading two terms, we use Stein’s lemma for vector-valued functions (e.g. see [27, Lemma 2.2]),
@@
7.1.2. Proof of $(\delta_{n},\eta_{n})$- localizability
@@
We verify $(\delta_{n},\eta_{n})$-localizability for tensor pca with $\tilde{u}(x)=\sqrt{n}(m(x),r_{\perp}^{2}(x))$. We note that verification will then follow immediately for the cases $u(x)=(\sqrt{n}m(x),r_{\perp}^{2}(x))$ and $u(x)=(m(x),r_{\perp}^{2}(x))$ or any centering of these statistics. We also note item $(1)$ of $(\delta_{n},\eta_{n})$-localizability has already been verified for the no-preconditioning case (e.g. see [5, Section 7]). We move on to $(2)$:
$\sup_{x\in u^{-1}(E_{k})}\|\nabla\tilde{\Phi}\|$ $=\sup_{x\in u^{-1}(E_{k})}\sqrt{n}\frac{\|\nabla\Phi(x)\|}{\sqrt{\mathbb{E}\|\nabla L(x,y)\|^{2}}}+O(1/\sqrt{n})$
@@
$\sup_{x\in u^{-1}(E_{k})}\|\nabla\tilde{\Phi}\|$ $\lesssim\sup_{x\in u^{-1}(E_{k})}\sqrt{n}\left[\frac{C_{k}[\|x\|^{2k-2}+\|x\|^{k-1}]}{\sqrt{n\|x\|^{2k-2}+\|\Phi(x)\|^{2}}}+O\left(\frac{1}{\sqrt{n}}\right)\right]\leq C_{k}$
@@
$\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\|\nabla\tilde{H}(x,y)\|^{8}$ $\lesssim n^{4}\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\left[\left\|\frac{\nabla L(x,y)}{\|\nabla L(x,y)\|}\right\|^{8}+\|\nabla\tilde{\Phi}(x)]\|^{8}+O(n^{-4})\right]$
@@
$\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\langle\nabla\tilde{H}(x,y),\nabla\tilde{u}_{i}(x)\rangle^{4}$ $=\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\langle\frac{\sqrt{n}\nabla L(x,y)}{\|\nabla L(x,y)\|}-\nabla\tilde{\Phi}(x),\nabla\tilde{u}_{i}(x)\rangle^{4}$
$\lesssim n^{2}\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\langle\frac{\nabla L(x,y)}{\|\nabla L(x,y)\|},\nabla\tilde{u}_{i}(x)\rangle^{4}$
$\leq n^{2}\sup_{x\in u^{-1}(E_{k})}\sqrt{\mathbb{E}\langle\nabla L(x,y),\nabla\tilde{u}_{i}(x)\rangle^{8}\mathbb{E}\frac{1}{\|\nabla L(x,y)\|^{8}}}$
Where in the second line, we use $(a+b)^{4}\lesssim a^{4}+b^{4}$ and Jensen’s Inequality to remove the $\nabla\tilde{\Phi}$ term at the cost of a constant factor. We now recall that $\nabla L(x,y)$ is n-dimensional Gaussian with mean $O(1)$ and covariance $\Sigma=I_{d}+\lambda xx^{t},\|\Sigma\|_{op}=O(1),\|\Sigma\|_{F}=\Theta(n)$. For both $\sqrt{n}m$ and $\sqrt{n}r^{2}$ we have that $\langle\nabla L(x,y),\nabla\tilde{u}_{i}(x)\rangle$ is Gaussian with variance bounded by $C_{k}n$ and by a technical lemma for Gaussians we have that $\mathbb{E}\frac{1}{\|\nabla L(x,y)\|^{8}}=O(1/n^{4})$ yielding:
$\max_{i}\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\langle\sqrt{n}\nabla\tilde{H}(x,y),\nabla\tilde{u}_{i}(x)\rangle^{4}=O(n^{2})$
@@
$\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\langle\nabla^{2}u_{i}(x),\nabla\tilde{H}(x,y)\otimes\nabla\tilde{H}(x,y)-\tilde{V}\,\rangle^{2}$
$=n\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\,\text{tr}([I-vv^{t}](\nabla\tilde{H}(x,y)\otimes\nabla\tilde{H}(x,y)-\tilde{V}))^{2}$
$\lesssim n\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\,\text{tr}(\nabla\tilde{H}(x,y)\otimes\nabla\tilde{H}(x,y)-\tilde{V})^{2}+\mathbb{E}\text{tr}(vv^{\top}[\nabla\tilde{H}(x,y)\otimes\nabla\tilde{H}(x,y)-\tilde{V}])^{2}$
We now bound the first term above noting that the second term above is bounded by (e.g. using the Von Neumann trace inequality) the first term multiplied by $\|vv^{t}\|_{\text{op}}=O(1)$:
$\text{tr}(\nabla\tilde{H}(x,y)\otimes\nabla$ $\tilde{H}(x,y)-\tilde{V})=\,\text{tr}\left(\left[\frac{\sqrt{n}\nabla L(x,y)}{\|\nabla L(x,y)\|}-\nabla\tilde{\Phi}(x)\right]^{\otimes 2}-\mathbb{E}\left[\frac{\sqrt{n}\nabla L(x,y)}{\|\nabla L(x,y)\|}-\nabla\tilde{\Phi}\right]^{\otimes 2}\right)$
$=\,\text{tr}\left(\frac{\sqrt{n}\nabla L(x,y)^{\otimes 2}}{\|\nabla L(x,y)\|^{2}}\right)-2\text{tr}\left(\frac{\sqrt{n}\nabla L(x,y)}{\|\nabla L(x,y)\|}\otimes\nabla\tilde{\Phi}(x)\right)+\text{tr}(\nabla\tilde{\Phi}(x)^{\otimes 2})$
$-\mathbb{E}[\text{tr}\left(\frac{\sqrt{n}\nabla L(x,y)^{\otimes 2}}{\|\nabla L(x,y)\|^{2}}\right)-2\text{tr}\left(\frac{\sqrt{n}\nabla L(x,y)}{\|\nabla L(x,y)\|}\otimes\nabla\tilde{\Phi}(x)\right)+\text{tr}(\nabla\tilde{\Phi}(x)^{\otimes 2})]$
$=2\,\text{tr}\left(\frac{\sqrt{n}\nabla L(x,y)}{\|\nabla L(x,y)\|}\otimes\nabla\tilde{\Phi}(x)\right)-2\mathbb{E}\text{tr}\left(\frac{\sqrt{n}\nabla L(x,y)}{\|\nabla L(x,y)\|}\otimes\nabla\tilde{\Phi}(x)\right)$
@@
$\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\text{tr}(\nabla\tilde{H}(x,y)\otimes\nabla\tilde{H}(x,y)-\tilde{V})^{2}$ $\lesssim\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\text{tr}\left(\frac{\sqrt{n}\nabla L(x,y)}{\|\nabla L(x,y)\|}\otimes\nabla\tilde{\Phi}(x)\right)^{2}$
$\leq n^{2}\sup_{x\in u^{-1}(E_{k})}\|\nabla\tilde{\Phi}(x)\|^{2}$
$\leq C_{k}\,n$
where in the first line we use the fact $(a-b)^{2}\lesssim a^{2}+b^{2}$ and Jensen’s inequality. In the second line we use Cauchy Schwartz to bound the trace (which is here equal to an inner product) and in the last line we use item (2) of $(\delta_{n},\eta_{n})$-localizability above. Which yields the desired:
@@
Following Theorem 2.3 and the discussion following it, we find that the dynamics of SGD-M follow those of online SGD up to the additional $\beta$-dependent constants. With this, our analysis immediately follows from [5, Section 7] where we take the mappings,
@@
for $\beta\in(0,1)$. For example, by [5, Proposition 3.1], the limiting generators for online SGD are given by
@@
The proof of Proposition 3.2 follows directly from the case of online SGD in [5] and Theorem 2.3. In particular the equivalence of SGD and SGD-M after appropriate time and step-size rescaling. Below we prove Proposition 3.3.
@@
Applying Theorem 2.3 in light of 2.1, we simply check asymptotic closability, noting that $(\delta_{n},\eta_{n})$-localizability was already verified for this setting in the proof of Proposition 3.1. We check asymptotic closability which also follows easily from the same proof. The drifts under the new coordinates are written as
@@
$\tilde{f}_{\tilde{u}_{1}}=\begin{cases}\tilde{u}_{2}^{k-1}\tilde{u}_{1}-\lambda\left(\dfrac{\tilde{u}_{1}}{\sqrt{\tilde{u}_{2}}}\right)^{k-1}&k=2\\ \tilde{u}_{2}^{k-1}\tilde{u}_{1}&k\geq 3\end{cases},\qquad\tilde{f}_{\tilde{u}_{2}}=2\tilde{u}_{2}^{(k+1)/2}$
@@ 8. Proofs for Single-Index Models
We start with SGD-M below. The limiting generators were first determined in [43, Theorem 3.1, Corollary 3.2]. We provide a slightly different derivation of the result for completeness. Refer to [43, Lemma 3.5] to show that the problem satisfies the $\delta$-localizability and closability assumptions as they also apply to SGD-M.
@@
We now prove the remainder of the proposition for SGD-U. We note that verifying $(\delta_{n},\eta_{n})$-localizability is a similar exercise as to the case of Tensor PCA in Section 7.1.2. We thus cover the key observations below. Again, we consider here $\tilde{u}(x)=\sqrt{n}u(x)=\sqrt{n}(m(x),r_{\perp}^{2}(x))$ which implies all cases of interest (i.e. $m$ and $r^{2}$ up to a rescaling of $\sqrt{n}$ with appropriate centering).
Proof for SGD-U $(\delta_{n},\eta_{n})$-localizability.
Item (1) is the same as the SGD-M case above. For item (2), the preconditioned gradient is:
$\eta\nabla L$ $=\frac{\sqrt{n}}{2|\Delta\sigma^{\prime}(s)|\|a\|}(2\Delta\sigma^{\prime}(s)a)$
$=\sqrt{n}\ \text{sgn}(\Delta\sigma^{\prime}(s))\left[\frac{a_{1}v+a_{2}u}{\|a\|}+\frac{w}{\|a\|}\right]$
Which implies that $\|\tilde{\Phi}(x)\|=O(1)$ given the expectation of each fraction above is $O(1/\sqrt{n})$. The above also trivially gives $\|\nabla\tilde{H}\|^{8}=O(\delta_{n}^{-4})$, concluding item (2). For item (3) we once again (as in the tensor PCA example) have:
$\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\langle\nabla\tilde{H}(x,y),\nabla\tilde{u}_{i}(x)\rangle^{4}$ $\leq n^{2}\sup_{x\in u^{-1}(E_{k})}\sqrt{\mathbb{E}\langle\nabla L(x,y),\sqrt{n}\nabla u_{i}(x)\rangle^{8}\mathbb{E}\frac{1}{\|\nabla L(x,y)\|^{8}}}$
We remind the reader that $\nabla L=2\Delta f^{\prime}(s)a$ and $\nabla u(x)=(v,2(x-mv))$, hence $\langle\nabla L(x,y),\sqrt{n}\nabla u_{i}(x)\rangle=O(\sqrt{n})$. Additionally, $\mathbb{E}\|\nabla L(x,y)\|^{-1}=O(1/\sqrt{n})$ which yields:
$\sup_{x\in u^{-1}(E_{k})}\mathbb{E}\langle\nabla\tilde{H}(x,y),\nabla u_{i}(x)\rangle^{4}=O(n^{2})=O(\delta^{-2})$
The last part of item (3) follows identically to the Tensor PCA example. ∎
Proof for SGD-U asymptotic closability.
@@
where $S^{2}=1$ almost surely. Since $a$ is an isotropic Gaussian vector, $\hat{a}$ is uniformly distributed on the sphere. It follows that $\mathbb{E}[\hat{a}\hat{a}^{\top}]=I/n$ (see, for instance [55, Proposition 3.3.8] for the construction). It follows then that the covariance matrix is: